OpenStudy (anonymous):

factor: x^4+2x^3+4x^2+8x

OpenStudy (anonymous):

x(x^2+4)(x+2)

OpenStudy (anonymous):

lol sorry not to turn you down but I have the answer key for this, and thats not the answer. i'm wondering how they got the answer. the answer is: (x^3+4x)(x+2)

OpenStudy (anonymous):

x^4+2x^3+4x^2+8x =x^4+4x^2+2x^3+8x =x^2(x^2+4)+2x(x^2+4) =(x^2+2x)(x^2+4) =x(x+2)(x^2+4) =x(x^2+4)(x+2) =(x^3+4x)(x+2)